A store sells bluegrass seed for $6 per pound and rye grass seed for $3 per pound. How much rye grass must be mixed with 100 pounds of bluegrass to obtain a blend that will sell for $5 per pound?
Answer by jorel1380(3719)

100(6)+m(3)=(100+m)(5)
600+3m=500+5m
100=2m
m=50
The store needs 50 lbs. of rye grass for their mix..